OpenStudy (anonymous):

Someone please help me with my homework! It deals with arithmetic operations using like and unlike radical terms and expressing the result in simplest form.

OpenStudy (anonymous):

I am on #10 on my homework, it says.. Darin simplified \[5\sqrt{\sqrt{5}}+2\sqrt{5}\] and got 15.7 Martha simplified the same expression and got 50. Use a calculator to determine who got the correct answer.

OpenStudy (hba):

Well you can decide that easily \[\sqrt{a}=a^{1/2}\]

OpenStudy (anonymous):

How did you get that?

OpenStudy (hba):

What ?

OpenStudy (anonymous):

How did you get your answer..?

OpenStudy (hba):

That is not the answer I am explaining you that anything in \[\sqrt{} \ becomes \ ^{1/2}\]

OpenStudy (hba):

Similarly you can change \[\sqrt{5}=5^{1/2}\]